OpenStudy (loser66):

Find \(lim_{x\rightarrow 0}(\dfrac{1}{x^4}-\dfrac{3x-x^3}{3arctan x})\) Please, help

ganeshie8 (ganeshie8):

\[\large \lim_{x\rightarrow 0}\dfrac{3x-x^3}{3\arctan x} \] you may start by showing the above limit is finite\(\), then the actual limit in question becomes \(\infty\) as \(\lim\limits_{x\to 0} \dfrac{1}{x^4} = \infty\)

OpenStudy (loser66):

Thank you, I got it.